OpenStudy (jennychan12):

Integrate. integral (x^1/3/x^2/3(1+x^1/3))dx

OpenStudy (jennychan12):

\[\int\limits (\frac{ x^{1\frac{ 1 }{ 3 }} }{ x^{\frac{ 2 }{ 3 }}(1+x^{\frac{ 1 }{ 3 }}) })dx\]

OpenStudy (jennychan12):

whoops that extra 1 on the top should be there. i set u = 1+x^1/3

OpenStudy (jennychan12):

oops, i meant that extra 1 shouldn't be there. du = 1/3x^-2/3 so the new integral should be \[\int\limits (\frac{ u+1 }{ u })du\] right?

OpenStudy (anonymous):

\[\int\frac{x^{\frac{1}{3}}}{x^{\frac{2}{3}}\left(1+x^{\frac{1}{3}}\right)}\;dx\\ \begin{align*}\text{Letting } u&=1+x^{\frac{1}{3}}\\ du&=\frac{1}{3}x^{-\frac{2}{3}}\;dx\end{align*}\\ 3\int\frac{u-1}{u}\;du\] Close, though!

OpenStudy (jennychan12):

yeah i realized i typed my response wrong. yeah the top should be u -1 and there should be a 3 in the front. after i deleted my old response i forgot.

OpenStudy (anonymous):

The integral is much simpler here; think you can manage?

OpenStudy (jennychan12):

yeah i think so. thanks :)